# Open X-Embodiment: Robotic Learning Datasets and RT-X Models

### v7

@@ I Introduction
00footnotetext: 1Allen Institute for AI; 2Arizona State University; 3California Institute of Technology; 4Carnegie Mellon University; 5Columbia University; 6EPFL; 7ETH Zürich; 8Flexiv Robotics; 9Georgia Institute of Technology; 10German Aerospace Center; 11Google DeepMind; 12Google Research; 13IO-AI TECH; 14Imperial College London; 15Intrinsic LLC; 16Istituto Italiano di Tecnologia; 17Korea Advanced Institute of Science & Technology; 18Max Planck Institute; 19Meta AI; 20Microsoft Research; 21Mila Quebec; 22NVIDIA; 23New York University; 24Princeton University; 25Queensland University of Technology; 26RIKEN; 27Shanghai Jiao Tong University; 28Stanford University; 29Technische Universität Darmstadt; 30The University of Texas at Austin; 31The University of Tokyo; 32Toyota Research Institute; 33University of California, Berkeley; 34University of California, Davis; 35University of California, San Diego; 36University of Edinburgh; 37University of Freiburg; 38University of Illinois Urbana-Champaign; 39University of Michigan; 40University of Montreal; 41University of Pennsylvania; 42University of Southern California; 43University of Technology, Nuremberg; 44University of Texas at Austin; 45University of Washington

### v8

@@ I Introduction
00footnotetext: 1Allen Institute for AI; 2Arizona State University; 3California Institute of Technology; 4Carnegie Mellon University; 5Columbia University; 6EPFL; 7ETH Zürich; 8Flexiv Robotics; 9Georgia Institute of Technology; 10German Aerospace Center; 11Google DeepMind; 12Google Research; 13IO-AI TECH; 14Imperial College London; 15Intrinsic LLC; 16Istituto Italiano di Tecnologia; 17Korea Advanced Institute of Science & Technology; 18Max Planck Institute; 19Meta AI; 20Microsoft Research; 21Mila Quebec; 22NVIDIA; 23New York University; 24Princeton University; 25Queensland University of Technology; 26RIKEN; 27Shanghai Jiao Tong University; 28Stanford University; 29Technische Universität Darmstadt; 30The University of Texas at Austin; 31The University of Tokyo; 32Toyota Research Institute; 33Tsinghua University; 34University of California, Berkeley; 35University of California, Davis; 36University of California, San Diego; 37University of Edinburgh; 38University of Freiburg; 39University of Illinois Urbana-Champaign; 40University of Michigan; 41University of Montreal; 42University of Pennsylvania; 43University of Southern California; 44University of Technology, Nuremberg; 45University of Texas at Austin; 46University of Washington